OpenStudy (anonymous):

an isotope is 4.5 x 10^9 years old. after a certain amount of time , it is now 1. 8 x 10^10 years old . In fraction form , what amount remained ? in other words , how much of 4.5 * 10^9 is 1.8 x 10^10 ?

Parth (parthkohli):

\(\LARGE \color{purple}{\rightarrow {4500000000 \over 18000000000} }\) Do it now?

OpenStudy (ash2326):

@princesskks did you understand?

Parth (parthkohli):

It's represented in scientific notation in the question.

OpenStudy (anonymous):

@ash2326 yes i do thanks @ParthKohli i got 0.25 . is that right ?

Parth (parthkohli):

Yes, you cancel out and get 45/180 = 1/4 = 0.25

OpenStudy (anonymous):

thanks !! @ParthKohli